MATH QUESTION~
A wheel rests against a floor and a vertical wall. A point P on the rim of the wheel is 2cm and 9cm from the floor and wall respectively Calculate the radius of the wheel.
Please help!
Thank you!

Answer by htmentor(1343)

If the wheel is resting on the floor and against the wall, the tangent lines are x=0 and y=0 and the points of tangency are (r,0) and (0,r).
The equations for the lines perpendicular to these tangent lines at these points are x=r and y=r and they must go through the center of the circle.
So the center of the circle is (r,r)
The standard form for a circle is (x-a)^2 + (y-b)^2 = r^2 where (a,b) is the center and r is the radius
Since the point P is 2 cm from the floor and 9 cm from the wall, the point is (9,2)
Inserting the point (9,2) in the equation for the circle gives
(9-r)^2 (2-r)^2 = r^2
Simplify and solve for r:
81 - 18r + r^2 + 4 - 4r + r^2 = r^2
r^2 - 22r + 85 = 0
This can be factored as
(r-17)(r-5) = 0
This gives two possible solutions, r=5 cm and r=17 cm
The graph below shows the two circles which share the point (9,2)